Circular No. 8718

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 HW_{51} (SIDING SPRING) An apparently asteroidal object discovered by R. H. McNaught in the

course of the Siding Spring project (discovery observation tabulated below) has been found to be cometary by A. Fitzsimmons, Department of Physics and Astronomy, Queen’s University, Belfast (together with S. C. Lowry and C. Snodgrass) on images taken with the 2.0-m Faulkes Telescope North on June 4.3 UT in poor seeing (2′′.0 FWHM), the 300-s R-band exposure showing a significantly flatter profile than exhibited by field stars, with a roughly symmetric coma extending 3′′; a slightly longer exposure on June 5.3 in 1′′.6 seeing shows a coma extending 3′′ from the central condensation (magnitudes within a 3′′-radius aperture were V = 17.77, V − R = +0.46).

2006 UT  α_{2000}  δ_{2000}  Mag.
Apr. 23.48710  9^{h}38^{m}29.^{s}94  −37^{o}16′47′′.3  17.5

Recent astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2006-L27.

T = 2006 Sept.29.4083 TT  ω = 359.^{o}9812
Ω = 228.1499 }2000.0
q = 2.265030 AU  i = 45.7966

COMETS C/2006 H3–H6 (SOHO) Additional Kreutz sungrazing comets have been found on SOHO web-

site images (cf. IAUC 8717; AA = A. Ambrus). C/2006 H3 was described by K. Battams as very faint, slightly diffuse, and elongated. C/2006 H4 appeared stellar in C3 images, reaching mag 6.1 on Apr. 21.988 at 10.6R_{⊙}; in C2 images, it showed a faint, thin tail that reached a length of 383′′ on Apr. 22.313. C/2006 H5 was very small, slightly diffuse, and also too faint for photometry. C/2006 H6, which was also found by H. Su, was small and stellar in C3 images, but elongated in C2 images, where it reached mag 6.4 on Apr. 28.603 at 6.6R_{⊙}.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 H3 Apr. 19.354 1^{h}56.^{m}3 +10^{o}16′ C2  KB 2006-K29
C/2006 H4  21.404 2 15.3 + 9 44 C3/2 AA 2006-K29
C/2006 H5 24.314 2 13.8 +11 51 C2 HS 2006-K29
C/2006 H6  28.321 2 32.9 +12 34 C3/2 JS 2006-L20
2006 June 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green